# Explain What You Mean: Intent Augmented Knowledge Graph Recommender Built With An LLM

### v2

@@
Abstract.
Interaction sparsity is a long-standing challenge in recommendation systems. Sparsity manifests in environments with disproportional cardinality of groupings of entities, such as users and products in an online marketplace. It is also found for newly introduced entities, described as the cold-start problem. Recent efforts to mitigate this issue either enrich the connectivity data by incorporating social networks or external knowledge graphs, or fine-tune LLMs into interaction augmenters or next-item recommenders. However, these techniques tend to be resource demanding, requiring high computational power. They also have several limitations, including data availability, low quality, or synthetic noise issues. In this work, we propose LLM-based Intent Knowledge Graph Recommender (IKGR), a novel framework that leverages retrieval-augmented generation and an encoding approach to construct and densify a knowledge graph. IKGR leverages latent user-item affinities from an interaction knowledge graph and further densifies it through mutual intent connectivity. This addresses sparsity issues and allows the model to make intent-grounded recommendations with an interpretable embedding translation layer. Through extensive experiments on real-world datasets, we demonstrate that IKGR overcomes knowledge gaps and achieves substantial gains over state-of-the-art baselines on both publicly available and our internal recommendation datasets.
Recommendation system, Large Language Model (LLM), Knowledge Graph, Retrieval Augmented Generation (RAG)
1. Introduction
Recommender systems fulfill a crucial role by helping users navigate through an ever-increasing catalog of online content and services to find what is most relevant for their personalized interests (Meng et al., 2023a; Wei et al., 2022). In enterprise applications, recommender algorithms are also widely deployed in internal search platforms or forums to help employees and developers ease the understanding of enterprise knowledge and the discovery of helpful articles, tips, or properties (Verma et al., 2015; Roy and Dutta, 2022).
A wide family of recommendation strategies has demonstrated success in modeling users’ historical interactions, such as collaborative filtering (CF) (Koren et al., 2021), graph-based methods (Pei et al., 2019; Liu et al., 2020; Zuo et al., 2024), and more recently recommenders based on Large Language Models (LLMs) (Kim et al., 2024; Bao et al., 2023; Wu et al., 2024). As these methods rely on implicit feedback from previous interactions, their performance tends to degrade for cold-starting new users or items with sparse interactions, hence it has always been a central topic in recommender systems to efficiently leverage user/item features and interactions to overcome sparse implicit feedback.
To address the cold start problem, researchers have proposed various approaches to augment the plain user-item interaction data with additional information. Depending on the source of added contextual information, these augmentation methods can be grouped into three categories: (1) enhancing user/item attributes with existing data, such as item attributes (Ying et al., 2023), user-generated content (Fan et al., 2019; Meng et al., 2023b), or multi-modal features (Yi et al., 2022); (2) enhancing linkages with existing data, such as social networks (Jamali and Ester, 2010) or external knowledge graphs (Wang et al., 2018), and (3) enhancing user/item links and attributes with LLMs, such as interaction augmentation (Wei et al., 2024). These approaches attempt to enrich entity attributes, increase interactions, or add contextual information to help the recommender learn better representations and provide improved recommendations. However, these augmentation methods can still face several challenges, which we categorize as follows.
i) Synthetic data noise caused by preference differences and knowledge gaps. Handling noisy data is a long-standing challenge faced by recommender systems trained on implicit feedback, where temporary or individualized user preferences can induce false positive or false negative samples. The noise issue can be further exacerbated by certain attempts to alleviate sparsity, such as the generation of synthetic user interactions with LLMs (Wei et al., 2024). LLM-generated preferences can introduce additional distribution skewness, potentially over-representing popular items, and failing to capture the long-tail distribution of real-world user preferences. To mitigate the additional data noise, careful calibration techniques need to be employed, e.g., loss calibration and training sample selection (Wei et al., 2024). Furthermore, LLMs may exhibit knowledge gaps, lacking necessary domain-specific context. In enterprise use cases especially, there is often a gap between private enterprise data and public corpora, leading to difficulty for both LLMs and traditional data augmentation approaches.
ii) Excessive resource budget. Another line of research has leveraged LLMs as an end-to-end function to perform the recommendation task (Wu et al., 2024; Deldjoo et al., 2024; Bao et al., 2023; Kim et al., 2024). However, fine-tuning the LLM to mitigate the knowledge gap tends to become resource intensive both in terms of the computation budget and developer efforts to calibrate the data and the model for each specific use case.
iii) Lack of fine-grained understanding. Among the LLM-based linking approaches, their predictions are mostly based on natural language representations rather than fact-based knowledge graph structures, hindering interpretability and targeted adjustments. For example, when recommending popular head items, it may not be clear if the recommendation is based on a genuine user affinity or simply the popularity of the item. A bias that is hard to correct without understanding the LLM’s underlying reasoning. Furthermore, as the number of real user interactions grows over time, the relative value of LLM-synthesized edges may diminish. Existing node enrichment methods use side information to augment node attributes with text-based features, but typically do not decompose them into graph structures, preventing the use of graph-based inference techniques (Ko et al., 2022; Liang et al., 2023).
@@ 1. Introduction
Figure 1. In enterprise search, the queries could contain special terminologies and acronyms, where traditional search engines or personalized rerankers fail to capture the real intents under such knowledge gap. IKGR addresses the challenge via injecting fine-grained understanding to textual features.
Additionally, these approaches may generate abbreviated textual descriptions, which can impair the model’s representations and subsequent recommendations.
In order to build a recommender that is good at handling the knowledge gap mentioned above, this work proposes to leverage the general reasoning ability of the LLM to augment the interactions. We take the the enterprise search as a primary case study as visualized in Figure 1, while keeping our approach applicable to more general scenarios. Our objective is to address the issue of noisy LLM generations under knowledge gap while maximally reducing the computation budget. We propose Intent Knowledge Graph Recommender (IKGR), an LLM-based pipeline, to explicitly augment node-level information. We build a knowledge graph that is naturally linked to the user-item interaction graph, as a way to enhance user/item node side information. Inspired by their impressive natural language understanding ability, we utilize LLMs to augment the interaction graph. We also simplify the task for the LLM to only boost the node-level understanding, so that more reliable results can be expected. In the proposed IKGR pipeline, the LLM is asked to perform a node-level retrieval task, which returns a list of entities that reflect users’ consumption intents. To overcome the knowledge gap, we carefully prepare an external knowledge base, explaining common abbreviations and terms. To reduce the computation budget, we design the framework as tuning-free, while retrieving the knowledge explanations from the knowledge base and feed to the LLM via in-context learning. We also design the backend method to be a lightweight Graph Neural Network (GNN) based on embedding translation, which naturally leverages the learned intent representation to inject structure prior into the graph.
As the main contribution of this work, we see two major advantages of using an LLM to augment the knowledge graph in the IKGR pipeline. Firstly, IKGR embraces the shift from an heuristic embedding learning structure to an LLM-based explicit knowledge extraction regime. It leverages the LLM to predict a new type of node: the interaction intent from a natural language perspective. Unlike previous approaches that rely on uninterpretable vector embeddings to capture multifaceted node attributes, IKGR recognizes that sub-node intents written in natural language format are often extractable from node features based on LLM. The natural language representation provide a more intuitive reflection of user preferences and interaction intents compared to indirect embeddings, and could be more flexible in terms of intent quantities. By incorporating LLMs, IKGR captures the richness and context of natural language, enhancing the accuracy and effectiveness of recommendations. Secondly, the low-quality and incomplete side information is enhanced by leveraging the extensive knowledge of LLMs gained through the pre-training stage, which allows for a tuning-free model that is both aware of vast real-world knowledge and understands user preferences and interaction intents. This integration prevents the dispersion of features across separate vector spaces and provides more accurate results.
  • We tackle the data sparsity challenge in recommendation systems, and propose IKGR, a knowledge graph augmentation approach with the LLM. IKGR leverages in-context learning with LLM to mitigate knowledge gap, and focuses on the node-level intent understanding and linkage to avoid introducing noise in synthesizing interactions.
  • The proposed IKGR is LLM-tuning-free, and is able to build knowledge graph inherently linked to the user-item interactions without the need for entity alignment. IKGR further leverages the learned intent nodes to inject structural prior in the embedding space with a knowledge graph translation layer.
  • Results on the enterprise search dataset demonstrate that the proposed IKGR effectively mitigate the knowledge gap, and outperforms the state-of-the-art baselines on search engine optimization reranking. Results on open-sourced datasets also showed the superiority of IKGR on the general recommendation scenario.
2. Related Works
Personalized Re-ranking Recommender System. Re-ranking in recommender systems aims to refine an initially ranked list to better capture user preferences and item relationships. Traditional ranking methods optimize global performance but often overlook the mutual influence between items and user-specific intent (Pei et al., 2019). Recent approaches address these limitations by integrating semantic tag information (Liang et al., 2010), multi-stage ranking optimization (Evnine et al., 2024), and item relationships (Liu et al., 2020). Transformer-based models effectively model global item interactions, while graph-based methods leverage item relationships for improved ranking (Pei et al., 2019; Liu et al., 2020). Furthermore, self-supervised learning enhances sequential recommendation by mitigating data sparsity issues (Zhou et al., 2020). These methods collectively demonstrate that incorporating user personalization, item dependencies, and efficient ranking strategies significantly enhances re-ranking effectiveness.
@@ 2. Related Works
Recommender Approaches Discovering Intent. Recent works explore intent modeling by leveraging external signals such as search queries and knowledge graphs. UDITSR (Zhang et al., 2024) jointly models search and recommendation by utilizing explicit search queries to infer implicit demand intents, while a dual-intent translation mechanism captures relationships between inherent intent, demand intent, and item interactions. Alternatively, knowledge graph-based methods, such as KGIN (Wang et al., 2021), refine intent discovery by representing user-item interactions through fine-grained relational paths, improving interpretability and recommendation quality. These approaches demonstrate that integrating external intent signals and structured relational modeling enhances intent-aware recommendations.
Recommender Approaches Addressing Data Sparsity. Recent works tackle the data sparsity issue through enhanced model architectures, graph-based techniques, and knowledge-aware methods. The Item History Model (IHM) improves cold-start item recommendations by directly injecting user-interaction data into the item tower and employing an inductive structure for real-time inference (Zuo et al., 2024). Cold Brew addresses sparsity in graph neural networks by distilling node representations, mitigating the impact of missing or noisy neighbors (Zheng et al., 2021). RippleNet further alleviates sparsity by propagating user preferences through knowledge graph relations, enriching item representations beyond collaborative signals (Wang et al., 2018). These approaches demonstrate that leveraging historical interactions, distilling structural knowledge, and integrating external information effectively mitigate data sparsity in recommendation systems.
@@
LLM-Based Interaction Augmentation for Recommender Systems. Gen-RecSys provides a comprehensive overview of generative models in recommendation, highlighting LLM-driven natural language understanding and multimodal integration (Deldjoo et al., 2024). LLMRec introduces graph augmentation strategies using LLMs to enrich interaction graphs, refine side information, and denoise implicit feedback, demonstrating performance gains across benchmark datasets (Wei et al., 2024). BLAIR further bridges language and recommendation by pretraining sentence embeddings on large-scale review data, improving item retrieval in complex natural language contexts (Hou et al., 2024). Additionally, BERT4Rec employs bidirectional self-attention to model user behavior sequences, overcoming limitations of traditional sequential models (Sun et al., 2019). These studies collectively showcase the potential of LLMs in augmenting interactions, refining representations, and enhancing recommendation performance.
Directly Leveraging LLM as the Recommender. A-LLMRec enhances LLM-based recommendation by leveraging embeddings from state-of-the-art collaborative filtering models, excelling in both cold and warm scenarios while maintaining efficiency and model-agnostic integration (Kim et al., 2024). CoRAL introduces collaborative retrieval-augmented prompting, addressing LLMs’ reliance on semantic information by incorporating user-item interactions through reinforcement learning-based retrieval policies, significantly improving long-tail recommendation (Wu et al., 2024). TALLRec proposes a tuning framework to align LLMs with recommendation-specific tasks, demonstrating strong generalization and efficiency even with limited training data (Bao et al., 2023). These approaches collectively highlight the potential of LLMs as standalone recommenders by addressing cold-start challenges, enhancing collaborative reasoning, and improving alignment with recommendation-specific objectives.
3. Preliminaries
@@
This work brings techniques from recommendation systems, knowledge graph-based GNNs, and Retrieval Augmented Generation (RAG) with LLMs. To help readers better digest our terminology and approaches, we leverage this section to present the problem formulation and commonly adopted techniques across these domains, and introduce our contributions in Section 4.
3.1. Graph Formulation Of Recommendation System
@@
(1) $\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}})=\mathcal{E}_{\text{dev}})$
@@ 3.2. Recommendation with Data Augmentation
(2) $\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}}\cup\mathcal{G}^{+})=\mathcal{E}_{\text{dev}})$
@@
Figure 2. IKGR’s graph augmentation steps to build intent knowledge graph with LLM. IKGR overcomes knowledge gap via in-context learning, avoids synthetic noise by focusing on the simple node-level intent entity retrieval task, and being light weight and tuning-free.
3.3. Knowledge Graph Convolution Layer

### v3

# Tuning-Free LLM Can Build A Strong Recommender Under Sparse Connectivity And Knowledge Gap Via Extracting Intent

@@
Abstract
Recent advances in recommendation with large language models (LLMs) often rely on either commonsense augmentation at the item-category level or implicit intent modeling on existing knowledge graphs. However, such approaches struggle to capture grounded user intents and to handle sparsity and cold-start scenarios. In this work, we present LLM-based Intent Knowledge Graph Recommender (IKGR), a novel framework that constructs an intent-centric knowledge graph where both users and items are explicitly linked to intent nodes extracted by a tuning-free, RAG-guided LLM pipeline. By grounding intents in external knowledge sources and user profiles, IKGR canonically represents what a user seeks and what an item satisfies as first-class entities. To alleviate sparsity, we further introduce a mutual-intent connectivity densification strategy, which shortens semantic paths between users and long-tail items without requiring cross-graph fusion. Finally, a lightweight GNN layer is employed on top of the intent-enhanced graph to produce recommendation signals with low latency. Extensive experiments on public and enterprise datasets demonstrate that IKGR consistently outperforms strong baselines, particularly on cold-start and long-tail slices, while remaining efficient through a fully offline LLM pipeline.
1 Introduction
Modern recommender systems are expected to reason over sparse and evolving interactions while serving highly personalized needs across vast catalogs. This challenge is acute in enterprise environments, where recommendations support internal search and knowledge discovery yet must cope with heterogeneous vocabularies, domain jargon, and long-tail content [1, 2]. Collaborative filtering and graph-based models have improved user–item representation learning [3, 4, 5, 6], while knowledge-aware recommenders further leverage structured relations for interpretability [7, 8]. Nevertheless, their effectiveness is bounded by incomplete coverage and weak connectivity, particularly for long-tail and cold-start cases [9].
To mitigate incomplete coverage and weak connectivity, a prominent direction uses LLMs to inject commonsense relations, typically complement or substitute links then fuses them with an existing item KG (CSRec) [10]. This line is attractive because (i) it regularizes sparse graphs with priors that are broadly valid (jackets complement sweaters; lenses complement cameras), (ii) it is offline-friendly (relations are generated once and reused), and (iii) it improves item-side coverage where merchant metadata is incomplete. However, three limitations recur in practice. (1) Granularity & intent mismatch. Category-level commonsense smooths the space but only loosely correlates with user-specific intent. Two users who click the same camera page may have very different intents (e.g., low-light astrophotography vs. lightweight travel kit). Category edges cannot capture these fine distinctions, which limits cold-start personalization. (2) Cross-graph fusion & alignment risk. CSRec-style pipelines typically align an LLM-augmented commonsense graph with an existing metadata/interaction graph. This requires ontology matching, entity resolution, and confidence thresholds. Errors here introduce structural noise that is hard to debug, and different domains require repeated re-alignment [8]. (3) Temporal & domain drift. Off-the-shelf commonsense tends to be domain-neutral and slow-moving. Enterprise or fast-evolving domains (e.g., internal tools, niche APIs) quickly deviate; commonsense edges may become stale or irrelevant without domain grounding [11, 12].
Another line remains within a pre-existing KG and learns latent mixtures over relations to explain interactions [13, 14, 15]. This method has several strengths: it’s entirely symbolic and structural, avoiding the need for LLMs; it uses relation paths to capture higher-order semantics; and it doesn’t require cross-graph fusion because it operates within a single KG. However, these pre-existing KG methods have several notable gaps. First, the intent remains latent and difficult to audit, as it’s represented as a vector or a mixture over relations rather than an explicit, human-readable node. This limits explainability and hinders downstream applications like intent analytics. Additionally, the model is bounded by the existing KG, meaning that if the base KG is sparse or doesn’t align with how users actually express themselves, the model struggles to capture missing semantics and often overfits to popular hubs, which hurts its performance on the long tail [9]. Finally, there is a weak tie to external knowledge, which means issues like polysemy and synonymy in relation labels (e.g., guide, how-to, playbook) and unstandardized domain jargon persist, as the model lacks explicit text grounding or retrieval capabilities [16].
A third family places the LLM directly in the loop — either to synthesize interactions (augment clicks/purchases) or to produce rankings end-to-end [17, 18, 19]. This approach offers benefits such as broader semantic coverage and simplified modeling with fewer bespoke modules. However, the trade-offs are substantial. Inference-time LLM calls introduce significant latency and cost, complicating service-level agreements and A/B testing at scale, especially for multi-stage recommenders [12]. Furthermore, synthetic interactions often lead to distribution shift and popularity bias, as they over-represent head items and reflect the LLM’s generic priors rather than the platform’s actual demand patterns; even with loss calibration, realistic long-tail fidelity is difficult to guarantee [19]. Finally, governance and reproducibility are challenging due to frequent model or LoRA updates that change behavior and varying privacy filtering across deployments, making it difficult to maintain offline-online consistency and respond to incidents in enterprise environments [11].
@@ 1 Introduction
Figure 1: In enterprise search, the queries could contain special terminologies and acronyms, where traditional search engines or personalized rerankers fail to capture the real intents under such knowledge gap. IKGR addresses the challenge via injecting fine-grained understanding to textual features.
Summarizing the pros and cons of various existing methods, we distill three core challenges to LLM and KG based recommenders.
(i) User intent extraction is an effective way to clarify and densify the KG, but this step needs to be more grounded. User profiles, queries, and enterprise documents are messy: synonyms, abbreviations, internal codenames, polysemy, and multi-lingual snippets all coexist. Extracting what a user seeks and what an item satisfies requires more than NER; it needs disambiguation (MLP = multilayer perceptron vs. marketing launch plan), normalization to a stable vocabulary, and aggregation across sources. Pure LLM prompts help but are prone to hallucination and label drift across batches. Without external knowledge grounding (retrieval of glossaries, wikis, policy pages) and schema-aware canonicalization, intent nodes will be noisy, redundant, and unmaintainable [16, 11]. This suggests the approach to elevate intent to first-class nodes extracted and normalized via RAG-guided LLMs.
(ii) The extracted intent needs to be more directly aligned with the existing user and items. The goal is to shorten semantic paths by adding edges that reflect shared or similar intents (user $\leftrightarrow$ intent, item $\leftrightarrow$ intent), thereby enabling information to flow even when user–item links are missing. This step is better achieved by leveraging the extracted intents to densify a single graph, rather than through a separate cross-graph alignment stage. If densification requires merging a separate LLM commonsense graph with an interaction/metadata graph, we re-introduce entity resolution and ontology mismatch—exactly the failure modes that cause silent errors in production [10, 8].
(iii) The system should be efficient, stable, and scalable for deployment. An efficient system should keep all heavy LLM work offline. That means handling batch extraction, incremental refresh, and backfill for new content/users; caching retrieval results; and defining compatible online components (e.g., a small GNN layer) [12]. To avoid synthetic-data drift [19], improvements should also come from structural connectivity and explicit intents, not from generating pseudo-interactions.
In response to these needs, we propose LLM-based Intent Knowledge Graph Recommender (IKGR), an intent-centric knowledge graph framework. First, we instantiate intent nodes and link users and items to them via a tuning-free, RAG-guided LLM extraction pipeline, grounded in external knowledge [16, 11]. Second, we introduce mutual-intent connectivity densification, which structurally improves graph connectivity for sparse and cold-start regimes without relying on cross-graph fusion [10, 18]. Finally, a lightweight GNN layer learns over the intent-enhanced graph to produce low-latency recommendations, fully decoupled from online LLM inference [17, 12]. Empirically, IKGR consistently outperforms strong baselines on public and enterprise datasets, especially on long-tail slices, while providing interpretable pathways via explicit intents. The contributions of IKGR can be summarized as follows.
  • IKGR introduces an intent-centric KG construction approach that turns user and item intents into first-class nodes. Using a tuning-free, RAG-guided LLM extractor, it canonically links users and items to intents with high precision.
  • IKGR involves an intent connectivity densification step, which shortens semantic paths between users and long-tail items, improving sparsity and cold-start performance without cross-graph fusion.
  • IKGR’s pipeline is offline and low-latency online by construction, and consistently outperforms state-of-the-art KG or LLM baselines on public and enterprise datasets, with notable improvements on long-tail and cold-start slices.
2 Related Works
LLM and KG based Recommenders. Recent responses to sparsity and cold start fall into four camps. (1) LLM-assisted item-side augmentation injects category/attribute-level commonsense (e.g., complement/substitute) and fuses it into an item KG (CSRec) [10]. This regularizes sparse graphs with broadly valid priors and is offline-friendly, yet it struggles to align with user-specific intent (two camera clicks can imply very different needs), requires brittle cross-graph ontology/entity alignment, and drifts in fast-moving or enterprise domains without domain grounding [8, 11, 12]. (2) KG-based implicit-intent modeling (KGIN) stays inside a single KG and explains interactions via latent mixtures over relations [13]. It avoids fusion and can exploit multi-hop paths, but the intent remains a non-auditable latent vector, performance is bounded by the coverage of the base KG (hurting the long tail), and weak ties to text/external knowledge leave synonymy/polysemy unresolved [9, 16]. (3) LLM-as-Recommender / synthetic interactions broaden semantic coverage or even replace ranking end-to-end [17, 18, 19], but introduce inference-time cost and tail latency, amplify popularity bias in generated data, and complicate governance and reproducibility under frequent model updates [12, 11]. (4) Text/multimodal node enrichment improves representations by attaching reviews/UGC/features to nodes, yet leaves knowledge ungraphized—no new edges/nodes for message passing or structural densification [20, 21]. Taken together, current trends either inject coarse item-side priors, keep intent latent and hard to use beyond ranking, or pay online-LLM costs and stability penalties.
@@ 2 Related Works
Recommender Approaches Discovering Intent. Recent works explore intent modeling by leveraging external signals such as search queries and knowledge graphs. UDITSR [22] jointly models search and recommendation by utilizing explicit search queries to infer implicit demand intents, while a dual-intent translation mechanism captures relationships between inherent intent, demand intent, and item interactions. Alternatively, knowledge graph-based methods, such as KGIN [23], refine intent discovery by representing user-item interactions through fine-grained relational paths, improving interpretability and recommendation quality. These approaches demonstrate that integrating external intent signals and structured relational modeling enhances intent-aware recommendations.
@@
LLM-Based Interaction Augmentation for Recommender Systems. Gen-RecSys provides a comprehensive overview of generative models in recommendation, highlighting LLM-driven natural language understanding and multimodal integration [24]. LLMRec introduces graph augmentation strategies using LLMs to enrich interaction graphs, refine side information, and denoise implicit feedback, demonstrating performance gains across benchmark datasets [19]. BLAIR further bridges language and recommendation by pretraining sentence embeddings on large-scale review data, improving item retrieval in complex natural language contexts [25]. Additionally, BERT4Rec employs bidirectional self-attention to model user behavior sequences, overcoming limitations of traditional sequential models [26]. These studies collectively showcase the potential of LLMs in augmenting interactions, refining representations, and enhancing recommendation performance.
More related works regarding data sparsity, LLM as recommender, and personalized re-rankers are detailed in Appendix C.
3 Preliminaries
@@
We leverage this section to present the problem formulation and commonly adopted techniques across these domains, and introduce our contributions in Section 4.
Figure 2: IKGR’s graph augmentation steps to build intent knowledge graph with LLM. IKGR overcomes knowledge gap via in-context learning, avoids synthetic noise by focusing on the simple node-level intent entity retrieval task, and being light weight and tuning-free.
3.1 Graph Formulation of Recommendation System
@@
$\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}})=\mathcal{E}_{\text{dev}})$ (1)
@@
$\Theta^{*}=\operatorname*{arg\max}\limits_{\Theta}p(g_{\Theta}(\{(u,i)\}_{\text{dev}};\mathcal{G}_{\text{train}}\cup\mathcal{G}^{+})=\mathcal{E}_{\text{dev}})$ (2)
@@
3.2 Knowledge Graph Convolution Layer